# MK-SGN: A Spiking Graph Convolutional Network with Multimodal Fusion and Knowledge Distillation for Skeleton-based Action Recognition

### v4

@@ Abstract
In recent years, multimodal Graph Convolutional Networks (GCNs) have achieved remarkable performance in skeleton-based action recognition. The reliance on high-energy-consuming continuous floating-point operations inherent in GCN-based methods poses significant challenges for deployment in energy-constrained, battery-powered edge devices. To address these limitations, MK-SGN, a Spiking Graph Convolutional Network with Multimodal Fusion and Knowledge Distillation, is proposed to leverage the energy efficiency of Spiking Neural Networks (SNNs) for skeleton-based action recognition for the first time. By integrating the energy-saving properties of SNNs with the graph representation capabilities of GCNs, MK-SGN achieves significant reductions in energy consumption while maintaining competitive recognition accuracy. Firstly, we formulate a Spiking Multimodal Fusion (SMF) module to effectively fuse multimodal skeleton data represented as spike-form features. Secondly, we propose the Self-Attention Spiking Graph Convolution (SA-SGC) module and the Spiking Temporal Convolution (STC) module, to capture spatial relationships and temporal dynamics of spike-form features. Finally, we propose an integrated knowledge distillation strategy to transfer information from the multimodal GCN to the SGN, incorporating both intermediate-layer distillation and soft-label distillation to enhance the performance of the SGN. MK-SGN exhibits substantial advantages, surpassing state-of-the-art GCN frameworks in energy efficiency and outperforming state-of-the-art SNN frameworks in recognition accuracy. The proposed method achieves a remarkable reduction in energy consumption, exceeding 98% compared to conventional GCN-based approaches. This research establishes a robust baseline for developing high-performance, energy-efficient SNN-based models for skeleton-based action recognition
@@ 1 Introduction
Human action recognition is a critical task with wide-ranging applications in human-computer interaction, identity verification, and emergency detection [1]. Initial approaches predominantly relied on RGB-based methods that analyzed visual information extracted from video frames[2]. While effective in certain contexts, such methods encountered significant challenges in complex environments characterized by cluttered backgrounds, inadequate lighting, and occlusions, limiting their robustness and generalization capabilities.
@@
Figure 1: Comparsion of (a) GCN and (b) SGN for skeleton-based Action Recognition. The former performs feature extraction and information propagation on nodes of the graph structure through graph convolutional layers, while the latter employs spike encoding to convert features into discrete spike-form, conducting feature extraction and information propagation in the time dimension of the spike-form feature.
@@
To address limitations in earlier approaches, research has shifted towards skeleton-based action recognition, which utilizes 3D skeleton data obtained from depth sensors [3]. By isolating human action dynamics from environmental disturbances, skeleton-based recognition enhances robustness and reliability in action recognition tasks. Early methods employed Convolutional Neural Networks (CNNs) [4, 5, 6], Recurrent Neural Networks (RNNs) [8, 7], and Long Short-Term Memory networks (LSTMs) [9, 34] to model spatio-temporal features. However, limitations in representing the intrinsic graph structure of skeleton data prompted the development of Graph Convolutional Networks (GCNs), which represent a significant advancement in the field [10].
@@
GCNs effectively model the spatial and temporal relationships within skeleton data, with the introduction of ST-GCN representing a key milestone [11]. The architecture employs spatial-temporal graph convolution to process skeletal data and laid the foundation for more advanced frameworks. 2S-AGCN improved performance by integrating a two-stream structure (Joint and Bone Streams) alongside adaptive graph convolution, enabling dynamic learning of skeletal topologies and refinement of temporal features [12]. Subsequent advancements aimed to balance efficiency and accuracy. Shift-GCN utilized a Channel-wise Shift Operation to simulate spatio-temporal information flow with minimal computational overhead [13], whereas CTR-GCN introduced channel-level topology optimization and attention mechanisms to enhance feature representation [14]. Building on foundational models, Info-GCN leverages information bottleneck theory for representation learning and extends the skeleton modality to six streams using the k-hop approach [15]. The development culminated in HULK, a multimodal large-scale model designed to unify diverse modalities within a robust framework. By harnessing its extensive capacity, HULK achieves state-of-the-art(SOTA) performance in precision and multimodal integration [16].
@@
However, such performance in traditional methods often relies on deep network architectures typical of artificial neural networks (ANNs), which are heavily dependent on computationally expensive floating-point operations. Such reliance results in considerable computational energy demands. For example, lightweight models like Shift-GCN and CTR-GCN consume 46 mJ and 36.25 mJ per action sample, respectively, while 2S-AGCN requires up to 164.68 mJ per sample (single-sample energy consumption calculated based on Equation 37). Additionally, the pursuit of improved accuracy necessitates the integration of multimodal data, which requires separate training and testing for each modality, followed by ensemble fusion. These approaches significantly increase both computational complexity and energy consumption, posing substantial challenges for deployment on battery-powered edge devices and other energy-constrained environments
@@
Building on advancements in SNNs and GCNs, we propose a novel Spiking Graph Network (SGN) framework that, for the first time, integrates the energy efficiency of SNNs with the graph representation capabilities of GCNs for skeleton-based action recognition. Our work revolutionizes skeleton-based action recognition by establishing the first framework based on SNNs, unveiling their transformative potential in addressing diverse real-world challenges, particularly in environments with strict energy constraints.
@@
    Transformation of Skeleton Data into Spike-based Representations: What methodologies can effectively transform skeleton data, characterized by complex graphical structures, into spike-based representations compatible with the computational principles of SNNs?
@@
    Multimodal Feature Fusion within SNNs: What advanced techniques can facilitate the fusion of multimodal features in SNNs, ensuring consistency, coherence, and resilience in feature representation?
@@
    Compatibility with Vanilla GCNs: How can spike-based representations be effectively integrated with vanilla GCNs to ensure seamless functionality and robust performance in action recognition tasks?
@@
    Optimization of Energy Efficiency and Accuracy: What strategies can be employed to optimize the trade-off between energy consumption and recognition accuracy, achieving a balance that meets the requirements of energy-constrained environments?
The identified challenges emphasize crucial areas for innovation, driving progress at the intersection of SNNs and GCNs for achieving both energy-efficient and highly accurate action recognition. To address these issues, we first propose a Skeleton Spiking Coding (SSC) module to convert skeleton data into event spikes and a Spike-based Multimodal Fusion (SMF) module based on mutual information to compute the correlation between multimodal data for early fusion. The SMF module fully leverages the complementary information from multimodal skeleton data while avoiding the heavy training and testing burden associated with ensemble methods. Building on these components, we further develop a novel SGN block, which integrates Self-Attention-Spiking Graph Convolutions (SA-SGC) and Spiking Temporal Convolutions (STC). The SA-SGC employs a spiking attention mechanism to enhance the representation of spiking features, thereby improving classification performance. Meanwhile, the STC extracts temporal features from spike-form data derived from skeleton sequences, capturing fine-grained temporal dependencies to complement the spatial information encoded by the SA-SGC. This integrated design bridges the gap between SGN and traditional GCN in feature extraction and information propagation. Additionally, a GCN-to-SGN knowledge distillation method is formulated, comprising inner-layer feature distillation and soft label distillation. High-precision GCN models guide the training of SGN through these complementary strategies. A Feature Translation Module (FTM) is employed to convert inner-layer features from multimodal GCN into spike-form features, enabling effective inner-layer feature distillation. Soft label distillation transfers the output distribution knowledge from GCN to SGN, further enhancing model performance. The dual-distillation approach improves accuracy without increasing model complexity. Leveraging the maturity of GCN architectures and pre-trained weights, the process imposes no additional training burden. Ultimately, we propose Spiking Graph Convolutional Network with Multimodal Fusion and Knowledge Distillation (MK-SGN), the first framework to leverage SNNs for addressing skeleton-based action recognition. Extensive experiments validate the effectiveness of the proposed approach, demonstrating its ability to achieve energy-efficient and accurate performance while bridging the gap between SNNs and GCNs in this domain.
@@
(2) An SMF module is formulated to fuse multimodal skeleton data in the form of spike-form features based on mutual information, effectively integrating complementary information while minimizing energy consumption and computational overhead. The novel SA-SGC and STC modules further enhance the representation of spatial relationships and temporal dynamics within the fused spike-form features.
(3) A knowledge distillation strategy is formulated to transfer knowledge from GCN to SGN, integrating both inner-layer feature distillation and soft label distillation for heterogeneous networks. To enable effective intermediate layer distillation, an FTM is designed to convert features from the intermediate layers of multimodal GCN into spike-form representations.
(4) Extensive experiments on three challenging datasets demonstrate the advantages of our work. The theoretical energy consumption of MK-SGN is 0.614 mJ per action sample. Compared to SOTA GCN-based methods, it reduces energy consumption by more than 98%. Moreover, we surpass the accuracy of SOTA SNN-based methods.
@@ 2.1 Skeleton-based Action Recognition
In recent years, deep learning has significantly advanced skeleton-based action recognition [3]. Early research on skeleton data primarily uses CNNs to extract spatial features, focusing on patterns such as joint positions and spatial relationships [4, 5, 6]. RNNs, including the more advanced variant LSTMs, model temporal dynamics, capturing sequential dependencies in movement patterns over time. Such methods address spatial and temporal dimensions separately, with LSTMs excelling at learning long-term dependencies in motion sequences[8, 7, 34, 35, 36]. However, CNN- and RNN-based approaches fail to adequately exploit the graph-like structure of skeleton data, limiting their capacity to represent joint dependencies and hierarchical relationships fully.
Figure 2: The overview of the proposed network architecture. The teacher network consists of a 10-layer GCN that processes four multimodal skeleton datasets independently, followed by Global Average Pooling (GAP) and Fully Connected (FC) layers to generate soft labels. The student network is a 6-layer SGN that incorporates the following key modules: (a) The Spiking Coding Module (SSC), which transforms raw skeleton data into spike-form representations, enabling compatibility with spiking neural computations. (b) The Spiking Self-Attention Spiking Graph Convolution (SA-SGC) module, formulated to model the spatio-temporal dependencies within spike-form features effectively by combining self-attention and spiking graph convolution. (c) The Spiking Temporal Convolution (STC) module, which captures temporal patterns and refines temporal dynamics within spike-form features. Additionally, the Spike-based Multimodal Fusion Module (SMF) optimizes cross-modal integration using mutual information, and the Feature Transformation Module (FTM) maps inter-layer GCN features into spike-form representations for effective alignment with the SNN domain. Knowledge distillation is achieved through the FTM and supervision from the soft labels produced by the teacher network.
The introduction of ST-GCN marked a breakthrough by representing skeletons as graphs, enabling the explicit modeling of spatio-temporal relationships through graph convolutions[11]. Since then, numerous studies have aimed to improve GCN performance, focusing on both spatial and temporal modeling. In the spatial domain, methods have primarily focused on capturing intrinsic joint correlations through techniques such as learnable adjacency matrices [41, 42, 43], attention mechanisms [37, 38, 39, 40], and other adaptive strategies, to enhance spatial feature representations by dynamically modeling relationships between joints, thereby overcoming the limitations of static graph structures. In the temporal domain, methods have focused on improving the temporal receptive fields to effectively capture both long-term and short-term dependencies, employing multi-stage strategies to model complex action dynamics and achieve finer-grained temporal representations[14, 44]. The integration of multimodal skeleton data, including joint, bone, motion, and additional dynamic features, has significantly advanced the field. By leveraging complementary modalities, methods such as Shift-GCN, CTR-GCN, and Info-GCN enhance recognition performance [13, 14, 15]. The development culminated in Hulk [16], a multimodal large-scale model that unified diverse modalities into a comprehensive framework, achieving SOTA results. Despite advancements in skeleton-based action recognition, current methods rely on ANN-based architectures, which require significant floating-point operations and lead to high energy consumption. Although effective, scalability and energy efficiency remain challenges. SNNs present a potential solution but have yet to be explored in this context, leaving a notable gap in the research.
@@ 2.2 Spiking Neural Networks
Unlike ANNs, SNNs transmit information through discrete spike sequences, a mechanism inspired by biological principles. This spike-driven mode closely mimics brain function, offering significant potential for energy efficiency [17]. Early research in SNNs focused on various spiking neuron models designed to convert continuous inputs into spike sequences. Notable models include the Integrate-and-Fire (IF) neuron [45], which simplifies the process by accumulating input signals until a threshold is reached, and the Hodgkin-Huxley (HH) model [46], which provides a biophysically detailed representation of neuronal activity based on ion channel dynamics. The Leaky Integrate-and-Fire (LIF) neuron is widely regarded for its simplicity and efficiency [47]. The dynamics and update of the membrane potential are described as:
@@
$V(t+\Delta t)=\begin{cases}V_{\text{reset}},&\text{if }V(t)\geq V_{\text{thresh}}\\ V(t)+\left(\frac{-\left(V(t)-V_{\text{rest}}\right)+I(t)}{\tau_{m}}\right)\Delta t,&\text{otherwise}\end{cases},$ (2)
@@
These advancements have led to the development of various spiking neural architectures, such as Spiking Graph Convolutional Networks [21, 22, 23], Spiking Transformers[27, 28, 29], Spiking Recurrent Neural Networks [24, 25, 26], and Spiking Convolutional Neural Networks [48, 49]. While existing research has primarily focused on traditional ANN tasks, significant progress has been made in spiking adaptations of Transformer models, particularly for RGB image processing. Studies in this area have demonstrated substantial energy savings without compromising accuracy[51]. Other research has focused on event-driven recognition, which is particularly well-suited for SNNs due to their ability to process sparse and asynchronous data [50]. Although SNNs have made significant progress in other areas, their application to skeleton-based action recognition remains underexplored. The event-driven nature of SNNs provides a promising approach to significantly improving energy efficiency, albeit with a certain trade-off in accuracy. Leveraging this potential could provide a viable solution to the challenge of balancing energy efficiency and accuracy in action recognition tasks.

### v5

@@ Abstract
In recent years, multimodal Graph Convolutional Networks (GCNs) have achieved remarkable performance in skeleton-based action recognition. The reliance on high-energy-consuming continuous floating-point operations inherent in GCN-based methods poses significant challenge for deployment in energy-constrained, battery-powered edge devices. To address the limitation, MK-SGN, a Spiking Graph Convolutional Network with Multimodal Fusion and Knowledge Distillation, is proposed to leverage the energy efficiency of Spiking Neural Networks (SNNs) for skeleton-based action recognition for the first time. By integrating the energy-saving properties of SNNs with the graph representation capabilities of GCNs, MK-SGN achieves significant reduction in energy consumption while maintaining competitive recognition accuracy. Firstly, we formulate a Spiking Multimodal Fusion (SMF) module to fuse the multimodal skeleton spike-form feature effectively. Secondly, we propose a Self-Attention Spiking Graph Convolution (SA-SGC) module and a Spiking Temporal Convolution (STC) module to capture spatial relationships and temporal dynamics of spike-form features, respectively. Finally, we propose an integrated knowledge distillation strategy to transfer information from the multimodal GCN to the SGN, incorporating inner-layer feature and soft-label distillation to enhance the SGN’s performance. MK-SGN exhibits substantial advantages, surpassing state-of-the-art GCN frameworks in energy efficiency and outperforming state-of-the-art SNN frameworks in recognition accuracy. The proposed method remarkably reduces energy consumption, with 98% less consumed than conventional GCN-based approaches. In general, this research establishes a strong baseline for developing high-performance, energy-efficient SNN-based models for skeleton-based action recognition.
@@ 1 Introduction
Human action recognition is a critical task with wide-ranging applications in human-computer interaction, identity verification, and emergency detection [10]. Initial approaches predominantly relied on RGB-based methods that analyzed visual information extracted from video frames[11]. While effective in certain contexts, such methods encountered significant challenges in complex environments characterized by cluttered backgrounds, inadequate illumination, and occlusions, limiting their robustness and generalization capabilities.
@@
Figure 1: Transformation from skeleton to spikes. The former GCN performs feature extraction and information propagation on nodes of the graph structure through graph convolutional layers, while the latter SGN employs spike encoding to convert features into discrete spike-form, conducting feature extraction and information propagation in the time dimension of the spike-form feature.
@@
Research has shifted towards skeleton-based action recognition to address limitations in these earlier approaches, which utilize 3D skeleton data obtained from depth sensors and pose estimation[12]. By isolating human action dynamics from environmental disturbances, skeleton-based recognition enhances robustness and reliability in action recognition tasks. Early methods employed Convolutional Neural Networks (CNNs) [13, 14, 15], Recurrent Neural Networks (RNNs) [17, 16], and Long Short-Term Memory networks (LSTMs) [18, 43] to model spatio-temporal features. However, limitations in representing the intrinsic structure of skeleton data prompted the development of Graph Convolutional Networks (GCNs), which represent a significant advancement in the field [19].
@@
GCNs effectively model the spatial and temporal relationships within skeleton data, with the introduction of ST-GCN representing a key milestone [20]. The architecture employs spatial-temporal graph convolution to process skeleton data, laying the foundation for more advanced frameworks. 2S-AGCN improves performance by integrating a two-stream structure (Joint and Bone Streams) alongside adaptive graph convolution, enabling dynamic learning of skeleton topologies and refinement of temporal features [21]. Subsequent advancements aim to balance efficiency and accuracy. Shift-GCN utilizes a channel-wise shift operation to simulate spatio-temporal information flow with lightweight computational overhead[22]. CTR-GCN introduces channel-level topology optimization and attention mechanisms to enhance feature representation [23]. Building on these, Info-GCN leverages information bottleneck theory for representation learning and extends the skeleton modality to six streams using the k-hop approach [24]. The development culminated in HULK, a multimodal large-scale model designed to unify diverse modalities within a robust framework. By harnessing its extensive capacity, HULK achieves state-of-the-art(SOTA) performance in precision and multimodal integration [25].
@@
However, such methods rely on deep network architectures typical of artificial neural networks (ANNs), which depend heavily on computationally expensive floating-point operations. Such reliance results in considerable computational energy demands. For example, lightweight models like Shift-GCN and CTR-GCN consume 46 mJ and 36.25 mJ per action sample, respectively, while 2S-AGCN requires up to 182.87 mJ per sample (single-sample energy consumption calculated based on Equation 36). Additionally, the pursuit of better performance necessitates the integration of multimodal data, which requires separate training and testing for each modality, followed by ensemble fusion. These characteristics of ANN-based methods lead to significant computational complexity and energy consumption, posing substantial challenges for deployment on battery-powered edge devices and other energy-constrained environments.
@@
Building on advancements in SNNs and GCNs, we propose a novel Spiking Graph Network (SGN) framework that, for the first time, integrates the energy efficiency of SNNs with the graph representation capabilities of GCNs for skeleton-based action recognition. Our work establishes the first framework leveraging SNNs for skeleton-based action recognition, significantly advancing this task’s potential in energy-constrained environments.
@@
    Transformation of Skeleton Data into Spike-based Representations: Effectively transforming skeleton data, characterized by complex graphical structures, into spike-based representations compatible with the computational principles of SNNs remains a crucial challenge.
@@
    Multimodal Feature Fusion within SNNs: Facilitating the fusion of multimodal features in SNNs requires advanced methods to ensure consistency, coherence, and resilience in feature representation.
@@
    Compatibility with Vanilla GCNs: Adapting the vanilla GCN framework to spike-based representations while ensuring robust performance in action recognition poses a significant research challenge.
@@
    Optimization of Energy Efficiency and Accuracy: Balancing energy consumption and recognition accuracy to meet the requirements of energy-constrained environments remains an essential goal.
To address these issues, we first propose a Skeleton Spiking Coding (SSC) module to convert skeleton data into event spikes and a Spike-based Multimodal Fusion (SMF) module based on mutual information to compute the correlation between multimodal data for early fusion. The SMF module fully leverages the complementary information from multimodal skeleton data while avoiding the heavy training and testing burden associated with ensemble methods. Based on these components, we propose a novel SGN block integrating a Self-Attention-Spiking Graph Convolutions (SA-SGC) module and a Spiking Temporal Convolutions (STC) module. The SA-SGC captures the spatial dependencies of skeleton sequences represented in spiking form through a graph topology and enhances them via a spiking attention mechanism, while the STC extracts temporal dependencies to complement the spatial information represented by the SA-SGC. This integrated design provides a unified framework that leverages the strengths of SGN and vanilla GCNs for effective feature extraction and information propagation. Additionally, a GCN-to-SGN knowledge distillation method is formulated, comprising inner-layer feature distillation and soft label distillation. High-precision GCN models guide the training of SGN through these complementary strategies. A Feature Translation Module (FTM) is formulated to convert inner-layer features from multimodal GCN into spike-form features, enabling effective inner-layer feature distillation. Soft label distillation transfers the output distribution knowledge from GCN to SGN, further enhancing model performance. The dual-distillation approach improves accuracy without increasing model complexity. Leveraging the maturity of GCN architectures and pre-trained weights, the process imposes no additional training burden. Ultimately, we propose Spiking Graph Convolutional Network with Multimodal Fusion and Knowledge Distillation (MK-SGN), the first framework to leverage SNNs for skeleton-based action recognition. Extensive experiments validate the effectiveness of the proposed approach, demonstrating its ability to achieve energy-efficient and accurate performance while bridging the gap between SNNs and GCNs in this domain.
@@
(2)An SMF module is formulated to fuse multimodal skeleton data in spike form based on mutual information, effectively integrating complementary information while reducing energy consumption and computational overhead. The novel SA-SGC and STC modules further enhance the representation of spatial relationships and temporal dynamics within the fused spike-form features, respectively.
(3)A knowledge distillation strategy is established to transfer knowledge from GCN to SGN, integrating inner-layer feature and soft-label distillation for heterogeneous networks. To enable effective inner-layer feature distillation, an FTM is employed to convert features from the intermediate layers of multimodal GCN into spike-form representations.
(4) Extensive experiments on three challenging datasets demonstrate the advantages of our work. The theoretical energy consumption of MK-SGN is 0.596 mJ per action sample. In general, Compared to SOTA GCN-based methods, with 98% less energy consumed than conventional GCN-based approaches. Moreover, we surpass the accuracy of SOTA SNN-based methods.
@@ 2.1 Skeleton-based Action Recognition
In recent years, deep learning has significantly advanced skeleton-based action recognition [12]. Early research primarily employed CNNs to extract spatial features, capturing joint positions and local spatial relationships [13, 14, 15]. RNNs, especially LSTMs, were then applied to model temporal dynamics, excelling at capturing long-term sequential dependencies in motion data [17, 16, 43, 44, 45]. While these approaches successfully addressed spatial and temporal dimensions separately, they failed to fully exploit the inherent graph structure of skeleton data, limiting their capacity to represent complex joint dependencies and hierarchical relationships.
The introduction of ST-GCN [20] marked a turning point by representing skeletons as graphs and applying graph convolutions to jointly capture spatial and temporal patterns. However, its reliance on a fixed adjacency matrix restricted adaptability to dynamic joint interactions. To overcome this limitation, subsequent studies proposed learnable adjacency matrices [50, 51, 52], which allow edges to be optimized for specific actions, though at the cost of higher model complexity and potential overfitting. Another line of work incorporates attention mechanisms [46, 47, 48, 49], which selectively emphasize informative joints and capture long-range dependencies, thereby enriching feature representations. Nevertheless, these methods often increase computational cost. Other adaptive strategies, such as multi-scale sampling or channel-mixing, further improve robustness and flexibility, but the balance among accuracy, efficiency, and generalizability remains an open challenge.
Beyond spatial modeling, researchers have also enhanced temporal modeling. MS-G3D [68] introduced multi-scale temporal modeling through dilated temporal windows and cross-spacetime graph convolutions, improving the capture of long-range temporal dependencies, while STA-GCN [53] used temporal adaptive graph convolutions to jointly model short- and long-term patterns.Together, these advances demonstrate the importance of refining both spatial and temporal receptive fields to improve overall recognition performance.
More recently, multimodal integration has become a key driver of progress. Approaches that incorporate joint, bone, motion, and other dynamic cues enrich feature diversity. With the development of multimodal foundation models, Hulk [25] unified heterogeneous skeleton modalities into a comprehensive framework, achieving state-of-the-art accuracy. BlockGCN [7] dynamically redefined topology awareness to better capture dependencies, offering accuracy comparable to Hulk while retaining lower computational cost. This contrast highlights the trade-off between large-scale multimodal integration and efficient graph-based modeling. The rise of large language models (LLMs) has also inspired novel directions. PURLS [8] leveraged GPT-3 to enhance zero-shot action recognition by generating semantic action descriptions, facilitating transfer to unseen classes. Neuron [1] advanced this line with evolving micro-prototypes for tighter alignment between skeleton data and semantic context. Similarly, CrossGLG [2] applied a global–local–global LLM-guided strategy to one-shot recognition, enriching feature abstraction across multiple levels. Although not directly targeting action recognition, M-R²ET [6] explored motion retargeting through decoupled skeleton and shape modeling, offering valuable insights into handling diverse skeleton structures and transfer learning.
Despite these advances, most existing methods remain grounded in conventional ANNs, which rely heavily on floating-point operations and thus incur high computational and energy costs. This limits their scalability and deployment in resource-constrained scenarios. SNNs, known for their event-driven and energy-efficient computation, present a promising alternative. However, their potential in skeleton-based action recognition has not yet been explored, leaving a critical research gap that our work aims to address.
@@ 2.2 Spiking Neural Networks
Unlike ANNs, SNNs transmit information through discrete spike sequences, a mechanism inspired by biological principles. This spike-driven mode closely mimics brain function, offering significant potential for energy efficiency [26]. Early research in SNNs focused on various spiking neuron models designed to convert continuous inputs into spike sequences. Notable models include the Integrate-and-Fire (IF) neuron [54], which simplifies the process by accumulating input signals until a threshold is reached. The Hodgkin-Huxley (HH) model [55] provides a biophysically detailed representation of neuronal activity based on ion channel dynamics. The Leaky Integrate-and-Fire (LIF) neuron is widely regarded for its simplicity and efficiency [56]. The dynamics and update of the membrane potential are described as follows:
@@
$V(t+\Delta t)=\begin{cases}V_{\text{reset}},&\text{\hskip-20.00003ptif }V(t)\geq V_{\text{thresh}}\\ V(t)+\left(\frac{-\left(V(t)-V_{\text{rest}}\right)+I(t)}{\tau_{m}}\right)\Delta t,&\text{otherwise}\end{cases},$ (2)
@@
These advancements have led to the development of various spiking neural architectures, such as Spiking Graph Convolutional Networks [30, 31, 32], which integrate Spiking Neural Networks into Graph Convolutional Networks to efficiently process dynamic graph data while reducing computational costs. In the domain of sequence processing, Spiking Transformers [36, 37, 38] leverage spiking self-attention mechanisms to replace traditional matrix multiplication with sparse operations, improving energy efficiency in tasks like RGB image processing. Meanwhile, Spiking Recurrent Neural Networks[33, 34, 35] combine the temporal dynamics of recurrent networks with spiking neuron models, offering superior performance in sequential tasks such as speech and gesture recognition, while significantly reducing power consumption. Lastly, Spiking Convolutional Neural Networks [57, 58] adapt spiking neurons to convolutional layers, enabling efficient processing of spatial correlations in image-based tasks, providing an energy-efficient alternative for real-time applications on neuromorphic hardware.
Studies in this area have demonstrated substantial energy savings without compromising accuracy[60]. Other research has focused on event-driven recognition, leveraging SNNs for their suitability in processing sparse and asynchronous data [59]. Although SNNs have made significant progress in other areas, their application to skeleton-based action recognition remains underexplored. The event-driven nature of SNNs provides a promising approach to significantly improving energy efficiency, albeit with an inevitable trade-off in accuracy. Leveraging this potential could provide a viable solution to balancing energy efficiency and accuracy in action recognition tasks.